# Magnetisation moment of a bounded 3D sample: asymptotic recovery from planar measurements on a large disk

### v4

@@ 3.2. Illustration of the basic idea of the derivation of the net moment estimates
Since the magnetisation distribution $\vec{\mathcal{M}}$ is compactly supported, the Fourier transforms ^$\hat{\mathcal{M}}_{j}\left(\mathbf{k},t_{3}\right)$, $j\in\left\{1,2,3\right\}$, are entire functions in $k_{1}$, $k_{2}\in\mathbb{C}$ for each $t_{3}\in Q_{3}$, according to the Paley-Wiener theory (see, e.g., [24, Thm 4.1]). In particular, power-series expansion of ^$\hat{\mathcal{M}}_{j}\left(k_{1},0,t_{3}\right)$, $j\in\left\{1,2,3\right\}$, about the origin $k_{1}=0$ of the complex plane (Re $\left(\text{Re }k_{1},\,\text{Im }k_{1}\right)$ gives
^$\hat{\mathcal{M}}_{j}\left(k_{1},0,t_{3}\right)=$$\hat{\mathcal{M}}_{j}\left(\mathbf{0},t_{3}\right)+\partial_{k_{1}}\hat{\mathcal{M}}_{j}\left(\mathbf{0},t_{3}\right)k_{1}+\dfrac{1}{2}\partial_{k_{1}}^{2}\hat{\mathcal{M}}_{j}\left(\mathbf{0},t_{3}\right)k_{1}^{2}+\mathcal{O}\left(\left|k_{1}\right|^{3}\right),\hskip 10.00002ptj\in\left\{1,2,3\right\}.$
@@
[…]
@@
Re $\text{Re }\hat{B}_{3}\left(k_{1},0,h\right)=$ $\pi m_{3}\left|k_{1}\right|-2\pi^{2}\left(\left\langle x_{1}M_{1}\right\rangle+\left\langle\left(h-x_{3}\right)M_{3}\right\rangle\right)\left|k_{1}\right|^{2}$
@@
Im $\text{Im }\hat{B}_{3}\left(k_{1},0,h\right)=$ $\pi m_{1}k_{1}-2\pi^{2}\left(\left\langle\left(h-x_{3}\right)M_{1}\right\rangle-\left\langle x_{1}M_{3}\right\rangle\right)k_{1}\left|k_{1}\right|$
@@
$\iint_{\mathbb{R}^{2}\backslash D_{A}}e^{2\pi ik_{1}x_{1}}B_{3}\left(\mathbf{x},h\right)\mathrm{d}^{2}x=$ $-\frac{m_{3}}{4\pi}\int_{A}^{\infty}\int_{0}^{2\pi}e^{2\pi ik_{1}r\cos\theta}\mathrm{d}\theta\frac{\mathrm{d}r}{r^{2}}$
@@
Re $\text{Re }\hat{B}_{3}\left(k_{1},0,h\right)=$ $\iint_{D_{A}}B_{3}\left(\mathbf{x},h\right)\mathrm{d}^{2}x-2\pi^{2}k_{1}^{2}\iint_{D_{A}}x_{1}^{2}B_{3}\left(\mathbf{x},h\right)\mathrm{d}^{2}x$
@@
Im $\text{Im }\hat{B}_{3}\left(k_{1},0,h\right)=$ $2\pi k_{1}\iint_{D_{A}}x_{1}B_{3}\left(\mathbf{x},h\right)\mathrm{d}^{2}x-\frac{4\pi^{3}}{3}k_{1}^{3}\iint_{D_{A}}x_{1}^{3}B_{3}\left(\mathbf{x},h\right)\mathrm{d}^{2}x$
@@
where we assumed that, for sufficiently small $\left|k_{1}\right|$, we have Im $\text{Im }\mathcal{R}_{A,k_{1}}=\mathcal{O}\left(1/A^{3}\right)$.
@@ 3.3. Rigorous analysis and higher-order asymptotic estimates
[…]
@@
We are going to pursue the idea outlined in the previous subsection. Namely, comparing a series expansion of (3.7) about $k_{1}=0$ with that of (3.10), we shall deduce a set of identities which relate magnetisation moments to the integrals of the measured data $B_{3}\left(\mathbf{x},h\right)$ on $D_{A}$. More precisely, using compactness of the support of the magnetisation $\vec{\mathcal{M}}$, it follows from (3.7) that Re $\text{Re }\hat{B}_{3}\left(k_{1},0,h\right)$ is a convergent series in powers of $\left|k_{1}\right|$ whereas Im $\text{Im }\hat{B}_{3}\left(k_{1},0,h\right)$ is a power series in $\left|k_{1}\right|$ multiplied by $k_{1}$. To facilitate the situation of matching the coefficients of different representations, we shall focus on the region $k_{1}>0$. Consequently, in what follows, evaluation at $k_{1}=0^{+}$ will mean the limiting value at $k_{1}=0$ taken from the positive semiaxis ($k_{1}>0$).
@@ 3.3.1. Tangential components of the net moment
$+\left.\partial_{k_{1}}^{2n+1}\left(\iint_{\mathbb{R}^{2}\backslash D_{A}}\sin\left(2\pi k_{1}x_{1}\right)B_{3}^{\text{asympt}}\left(\mathbf{x},h\right)\mathrm{d}^{2}x\right)\right|_{k_{1}=0^{+}}$
$+\left.\partial_{k_{1}}^{2n+1}\left(\iint_{\mathbb{R}^{2}\backslash D_{A}}\sin\left(2\pi k_{1}x_{1}\right)\left[B_{3}\left(\mathbf{x},h\right)-B_{3}^{\text{asympt}}\left(\mathbf{x},h\right)\right]\mathrm{d}^{2}x\right)\right|_{k_{1}=0^{+}}.$
@@
(3.42) $\mathcal{I}^{\sin}\equiv\mathcal{I}^{\sin}\left(k_{1},A\right):=\iint_{\mathbb{R}^{2}\backslash D_{A}}\sin\left(2\pi k_{1}x_{1}\right)B_{3}^{\text{asympt}}\left(\mathbf{x},h\right)\mathrm{d}^{2}x,$
@@
(3.43) $\mathcal{R}_{2n+1}^{\sin}\equiv\mathcal{R}_{2n+1}^{\sin}\left(A\right):=\left.\partial_{k_{1}}^{2n+1}\left(\iint_{\mathbb{R}^{2}\backslash D_{A}}\sin\left(2\pi k_{1}x_{1}\right)\left[B_{3}\left(\mathbf{x},h\right)-B_{3}^{\text{asympt}}\left(\mathbf{x},h\right)\right]\mathrm{d}^{2}x\right)\right|_{k_{1}=0^{+}},\hskip 10.00002ptn\in\mathbb{N}_{0},$
@@
Note that, in (3.49)–(3.50), we employed integration by parts twice using the asymptotic behavior of $J_{1}$ given in (A.6).
@@
(3.61) $-\frac{4\pi^{3}}{3}\iint_{D_{A}}x_{1}^{3}B_{3}\left(\mathbf{x},h\right)\mathrm{d}^{2}x+\pi^{4}A^{2}\left(\frac{c_{1,0}}{A}-\frac{5c_{3,0}+c_{1,2}}{6A^{3}}\right)+\frac{1}{6}\mathcal{R}_{3}^{\sin}=d_{3},$
@@
(3.63) $-\frac{\left(2\pi\right)^{7}}{5040}\iint_{D_{A}}x_{1}^{7}B_{3}\left(\mathbf{x},h\right)\mathrm{d}^{2}x+\frac{\left(2\pi\right)^{8}A^{6}}{92160}\left(\frac{c_{1,0}}{A}+\frac{9c_{3,0}+c_{1,2}}{6A^{3}}\right)+\frac{1}{7!}\mathcal{R}_{7}^{\sin}=d_{7},$
@@
(3.65) $-\frac{\left(2\pi\right)^{11}}{39916800}\iint_{D_{A}}x_{1}^{11}B_{3}\left(\mathbf{x},h\right)\mathrm{d}^{2}x+\frac{\left(2\pi\right)^{12}A^{10}}{17340825600}\left(\frac{98c_{1,0}}{9A}+\frac{13c_{3,0}+c_{1,2}}{A^{3}}\right)+\frac{1}{11!}\mathcal{R}_{11}^{\sin}=d_{11}.$
@@
[…]
@@
(3.68) $\mathcal{R}_{2n+1}^{\sin}=$ $\left.\partial_{k_{1}}^{2n+1}\left(\iint_{\mathbb{R}^{2}\backslash D_{A}}\sin\left(2\pi k_{1}x_{1}\right)L_{N}\left(\mathbf{x}\right)\mathrm{d}^{2}x\right)\right|_{k_{1}=0^{+}}$
$+\left.\partial_{k_{1}}^{2n+1}\left(\iint_{\mathbb{R}^{2}\backslash D_{A}}\sin\left(2\pi k_{1}x_{1}\right)\left[B_{3}\left(\mathbf{x},h\right)-B_{3}^{\text{asympt}}\left(\mathbf{x},h\right)-L_{N}\left(\mathbf{x}\right)\right]\mathrm{d}^{2}x\right)\right|_{k_{1}=0^{+}}.$
@@
and such a bound is possible due to the remainder estimate $\mathcal{O}\left(1/\left|\mathbf{x}\right|^{N+4}\right)$ in (3.67). Here, in the third line of the estimates, we used the fact that the integral in $r$ converges for $N>2n-1$, $0\leq n\leq 5$, which is true for $N>9$. For such $N$, the obtained estimate of order $\mathcal{O}\left(1/A^{N-2n}\right)$ is clearly even better than was aimed for (recall (3.66)).
@@
[…]
@@
where, in passing the differential operator $\partial_{k_{1}}^{2p+2}$ under the integral sign, we took into account the asymptotic behavior at infinity of $J_{1}$ given by (A.6) and, on the second line, employed the following identity valid for $k_{1}>0$:
@@
[…]
@@ 3.3.2. Normal component of the net moment
(3.96) $\mathcal{I}^{\cos}\equiv\mathcal{I}^{\cos}\left(k_{1},A\right):=\iint_{\mathbb{R}^{2}\backslash D_{A}}\cos\left(2\pi k_{1}x_{1}\right)B_{3}^{\text{asympt}}\left(\mathbf{x},h\right)\mathrm{d}^{2}x,$
@@
(3.97) $\mathcal{R}_{2n}^{\cos}\equiv\mathcal{R}_{2n}^{\cos}\left(A\right):=\left.\partial_{k_{1}}^{2n}\left(\iint_{\mathbb{R}^{2}\backslash D_{A}}\cos\left(2\pi k_{1}x_{1}\right)\left[B_{3}\left(\mathbf{x},h\right)-B_{3}^{\text{asympt}}\left(\mathbf{x},h\right)\right]\mathrm{d}^{2}x\right)\right|_{k_{1}=0^{+}},\hskip 10.00002ptn\in\mathbb{N}_{0},$
@@
[…]
@@
(3.115) $-\frac{\left(2\pi\right)^{2}}{2}\iint_{D_{A}}x_{1}^{2}B_{3}\left(\mathbf{x},h\right)\mathrm{d}^{2}x+\frac{\left(2\pi\right)^{3}}{4}A^{2}\left(\frac{c_{0,0}}{A}-\frac{131c_{2,0}+49c_{0,2}}{180A^{3}}\right)+\frac{1}{2}\mathcal{R}_{2}^{\cos}=d_{2},$
@@
(3.117) $-\frac{\left(2\pi\right)^{6}}{720}\iint_{D_{A}}x_{1}^{6}B_{3}\left(\mathbf{x},h\right)\mathrm{d}^{2}x+\frac{1}{11520}\left(2\pi\right)^{7}A^{6}\left(\frac{c_{0,0}}{A}+\frac{109c_{2,0}+11c_{0,2}}{72A^{3}}\right)+\frac{1}{6!}\mathcal{R}_{6}^{\cos}=d_{6},$
@@
(3.119) $-\frac{\left(2\pi\right)^{10}}{3628800}\iint_{D_{A}}x_{1}^{10}B_{3}\left(\mathbf{x},h\right)\mathrm{d}^{2}x+\frac{1}{132710400}\left(2\pi\right)^{11}A^{10}\left(\frac{c_{0,0}}{A}+\frac{523c_{2,0}+17c_{0,2}}{420A^{3}}\right)+\frac{1}{10!}\mathcal{R}_{10}^{\cos}=d_{10}.$
@@
[…]
@@
[…]
@@
[…]

### v5

@@ 3.2. Illustration of the basic idea of the derivation of the net moment estimates
Since the magnetisation distribution $\vec{\mathcal{M}}$ is compactly supported, the Fourier transforms $\hat{\mathcal{M}}_{j}\left(\mathbf{k},t_{3}\right)$, $j\in\left\{1,2,3\right\}$, are entire functions in $k_{1}$, $k_{2}\in\mathbb{C}$ for each $t_{3}\in Q_{3}$, according to the Paley-Wiener theory (see, e.g., [25, Thm 4.1]). In particular, power-series expansion of $\hat{\mathcal{M}}_{j}\left(k_{1},0,t_{3}\right)$, $j\in\left\{1,2,3\right\}$, about the origin $k_{1}=0$ of the complex plane $\left(\text{Re }k_{1},\,\text{Im }k_{1}\right)$ gives
$\hat{\mathcal{M}}_{j}\left(k_{1},0,t_{3}\right)=$ $\hat{\mathcal{M}}_{j}\left(\mathbf{0},t_{3}\right)+\partial_{k_{1}}\hat{\mathcal{M}}_{j}\left(\mathbf{0},t_{3}\right)k_{1}+\dfrac{1}{2}\partial_{k_{1}}^{2}\hat{\mathcal{M}}_{j}\left(\mathbf{0},t_{3}\right)k_{1}^{2}+\mathcal{O}\left(\left|k_{1}\right|^{3}\right),\hskip 10.00002ptj\in\left\{1,2,3\right\}.$
@@
$m_{j}=\int_{Q_{3}}\hat{\mathcal{M}}_{j}\left(\mathbf{0},t_{3}\right)dt_{3},\hskip 10.00002ptj\in\left\{1,2,3\right\},$
@@
$\text{Re }\hat{B}_{3}\left(k_{1},0,h\right)=$ $\pi m_{3}\left|k_{1}\right|-2\pi^{2}\left(\left\langle x_{1}M_{1}\right\rangle+\left\langle\left(h-x_{3}\right)M_{3}\right\rangle\right)\left|k_{1}\right|^{2}$
@@
$\text{Im }\hat{B}_{3}\left(k_{1},0,h\right)=$ $\pi m_{1}k_{1}-2\pi^{2}\left(\left\langle\left(h-x_{3}\right)M_{1}\right\rangle-\left\langle x_{1}M_{3}\right\rangle\right)k_{1}\left|k_{1}\right|$
@@
[…]
@@
$\text{Re }\hat{B}_{3}\left(k_{1},0,h\right)=$ $\iint_{D_{A}}B_{3}\left(\mathbf{x},h\right)\mathrm{d}^{2}x-2\pi^{2}k_{1}^{2}\iint_{D_{A}}x_{1}^{2}B_{3}\left(\mathbf{x},h\right)\mathrm{d}^{2}x$
@@
$\text{Im }\hat{B}_{3}\left(k_{1},0,h\right)=$ $2\pi k_{1}\iint_{D_{A}}x_{1}B_{3}\left(\mathbf{x},h\right)\mathrm{d}^{2}x-\frac{4\pi^{3}}{3}k_{1}^{3}\iint_{D_{A}}x_{1}^{3}B_{3}\left(\mathbf{x},h\right)\mathrm{d}^{2}x$
@@
where we assumed that, for sufficiently small $\left|k_{1}\right|$, we have $\text{Im }\mathcal{R}_{A,k_{1}}=\mathcal{O}\left(1/A^{3}\right)$.
@@ 3.3. Rigorous analysis and higher-order asymptotic estimates
(3.27) $c_{1,0}:=\frac{3}{4\pi}\left[\left\langle\left(h-x_{3}\right)M_{1}\right\rangle-\left\langle x_{1}M_{3}\right\rangle\right],\hskip 10.00002pt\hskip 10.00002ptc_{0,1}:=\frac{3}{4\pi}\left[\left\langle\left(h-x_{3}\right)M_{2}\right\rangle-\left\langle x_{2}M_{3}\right\rangle\right],$
@@
We are going to pursue the idea outlined in the previous subsection. Namely, comparing a series expansion of (3.7) about $k_{1}=0$ with that of (3.10), we shall deduce a set of identities which relate magnetisation moments to the integrals of the measured data $B_{3}\left(\mathbf{x},h\right)$ on $D_{A}$. More precisely, using compactness of the support of the magnetisation $\vec{\mathcal{M}}$, it follows from (3.7) that $\text{Re }\hat{B}_{3}\left(k_{1},0,h\right)$ is a convergent series in powers of $\left|k_{1}\right|$ whereas $\text{Im }\hat{B}_{3}\left(k_{1},0,h\right)$ is a power series in $\left|k_{1}\right|$ multiplied by $k_{1}$. To facilitate the situation of matching the coefficients of different representations, we shall focus on the region $k_{1}>0$. Consequently, in what follows, evaluation at $k_{1}=0^{+}$ will mean the limiting value at $k_{1}=0$ taken from the positive semiaxis ($k_{1}>0$).
@@ 3.3.1. Tangential components of the net moment
[…]
@@
[…]
@@
[…]
@@
Note that, in (3.49)–(3.50), we employed integration by parts twice using the asymptotic behaviour of $J_{1}$ given in (A.6).
@@
(3.61) $\frac{4\pi^{3}}{3}\iint_{D_{A}}x_{1}^{3}B_{3}\left(\mathbf{x},h\right)\mathrm{d}^{2}x+\pi^{4}A^{2}\left(\frac{c_{1,0}}{A}-\frac{5c_{3,0}+c_{1,2}}{6A^{3}}\right)+\frac{1}{6}\mathcal{R}_{3}^{\sin}=d_{3},$
@@
(3.63) $\frac{\left(2\pi\right)^{7}}{5040}\iint_{D_{A}}x_{1}^{7}B_{3}\left(\mathbf{x},h\right)\mathrm{d}^{2}x+\frac{\left(2\pi\right)^{8}A^{6}}{92160}\left(\frac{c_{1,0}}{A}+\frac{9c_{3,0}+c_{1,2}}{6A^{3}}\right)+\frac{1}{7!}\mathcal{R}_{7}^{\sin}=d_{7},$
@@
(3.65) $\frac{\left(2\pi\right)^{11}}{39916800}\iint_{D_{A}}x_{1}^{11}B_{3}\left(\mathbf{x},h\right)\mathrm{d}^{2}x+\frac{\left(2\pi\right)^{12}A^{10}}{17340825600}\left(\frac{98c_{1,0}}{9A}+\frac{13c_{3,0}+c_{1,2}}{A^{3}}\right)+\frac{1}{11!}\mathcal{R}_{11}^{\sin}=d_{11}.$
@@
(3.66) $\mathcal{R}_{2n+1}^{\sin}=\mathcal{O}\left(\frac{1}{A^{5-2n}}\right),\hskip 10.00002pt\hskip 10.00002pt0\leq n\leq 5.$
@@
[…]
@@
and such a bound is possible due to the remainder estimate $\mathcal{O}\left(1/\left|\mathbf{x}\right|^{N+4}\right)$ in (3.67). Here, in the third line of the estimates, we used the fact that the integral in $r$ converges for N>2n1N>2n-1, $0\leq n\leq 5$, which is true for $N>9$. For such $N$, the obtained estimate of order $\mathcal{O}\left(1/A^{N-2n}\right)$ is clearly even better than was aimed for (recall (3.66)).
@@
(3.70) $\mathcal{S}_{p,j}:=$ $\int_{A}^{\infty}\int_{0}^{2\pi}\sin\left(2\pi k_{1}r\cos\theta\right)\cos^{2j+1}\theta\,\mathrm{d}\theta\frac{\mathrm{d}r}{r^{2p+3}},\hskip 10.00002pt\hskip 10.00002pt0\leq j\leq p,\hskip 10.00002ptp\geq 2,$
@@
where, in passing the differential operator $\partial_{k_{1}}^{2p+2}$ under the integral sign, we took into account the asymptotic behaviour at infinity of $J_{1}$ given by (A.6) and, on the second line, employed the following identity valid for $k_{1}>0$:
@@
(3.84) $\frac{1}{2}\left(\mathcal{T}_{5}+\mathcal{T}_{9}\right)=\mathcal{T}_{7},\hskip 10.00002pt\hskip 10.00002pt\frac{1}{2}\left(\mathcal{T}_{7}+\mathcal{T}_{11}\right)=\mathcal{T}_{9}.$
@@ 3.3.2. Normal component of the net moment
[…]
@@
[…]
@@
(3.99) $\mathcal{I}_{1}^{\cos}:=\int_{A}^{\infty}\int_{0}^{2\pi}\cos\left(2\pi k_{1}r\cos\theta\right)d\theta\frac{\mathrm{d}r}{r^{2}},$
@@
(3.115) $\frac{\left(2\pi\right)^{2}}{2}\iint_{D_{A}}x_{1}^{2}B_{3}\left(\mathbf{x},h\right)\mathrm{d}^{2}x+\frac{\left(2\pi\right)^{3}}{4}A^{2}\left(\frac{c_{0,0}}{A}-\frac{131c_{2,0}+49c_{0,2}}{180A^{3}}\right)+\frac{1}{2}\mathcal{R}_{2}^{\cos}=d_{2},$
@@
(3.117) $\frac{\left(2\pi\right)^{6}}{720}\iint_{D_{A}}x_{1}^{6}B_{3}\left(\mathbf{x},h\right)\mathrm{d}^{2}x+\frac{1}{11520}\left(2\pi\right)^{7}A^{6}\left(\frac{c_{0,0}}{A}+\frac{109c_{2,0}+11c_{0,2}}{72A^{3}}\right)+\frac{1}{6!}\mathcal{R}_{6}^{\cos}=d_{6},$
@@
(3.119) $\frac{\left(2\pi\right)^{10}}{3628800}\iint_{D_{A}}x_{1}^{10}B_{3}\left(\mathbf{x},h\right)\mathrm{d}^{2}x+\frac{1}{132710400}\left(2\pi\right)^{11}A^{10}\left(\frac{c_{0,0}}{A}+\frac{523c_{2,0}+17c_{0,2}}{420A^{3}}\right)+\frac{1}{10!}\mathcal{R}_{10}^{\cos}=d_{10}.$
@@
(3.120) $\mathcal{R}_{2n}^{\cos}=\mathcal{O}\left(\frac{1}{A^{5-2n}}\right),\hskip 10.00002pt\hskip 10.00002pt0\leq n\leq 5.$
@@
(3.122) $\int_{A}^{\infty}\int_{0}^{2\pi}\cos\left(2\pi k_{1}r\cos\theta\right)\cos^{2\left(p-l\right)}\theta\,\,\cos^{2l}\theta\mathrm{d}\theta\frac{\mathrm{d}r}{r^{2\left(p+1\right)}},\hskip 10.00002pt\hskip 10.00002pt0\leq l\leq p,\hskip 10.00002ptp\geq 2,$
@@
(3.138) $\mathcal{T}_{0}=\mathcal{T}_{4}-2\mathcal{T}_{6},\hskip 10.00002pt\hskip 10.00002pt\mathcal{T}_{0}=4\mathcal{T}_{6}-25\mathcal{T}_{8}.$